OpenStudy (anonymous):

rewirte 19/5 as a mixed number 4 2/5 3 4/5 5 3/4 4 1/5

OpenStudy (anonymous):

B. 3 4/5

OpenStudy (anonymous):

Divide..

OpenStudy (anonymous):

Cuz 3*5=15 + 4 = 19 and u keep the denominator so u would get 19/5

OpenStudy (anonymous):

The number on top will be the whole number, the number at the bottom will be the numerator, and the divisor will be the numerator.

OpenStudy (anonymous):

So we have 3 4/5.

OpenStudy (anonymous):

or Divide the numerator by the denominator. For example, suppose you want to write the improper fraction 19/5 as a mixed number. First, divide 19 by 5: 19 ÷ 5 = 3 r 4 Write the mixed number in this way: The quotient (answer) is the whole-number part (3). The remainder of the improper fraction is the numerator (4). The denominator of the improper fraction is the denominator (5). So in this example, you would write the mixed number as follows: This method works for all improper fractions. And if you start with a reduced fraction, you don’t have to reduce your answer.

OpenStudy (anonymous):

Sorry about the ??? marks it always does that for some reason